"Word Problems Involving Products and Factoring"
Find two consecutive positive numbers such that the product of the sum and difference of the numbers plus 8 is the sum of their squares. can you solve this pls=( tyty so much

Answer by MRperkins(300)

Two consecutive positive numbers can be represented by n and n+1
the product of the sum and difference of the two numbers:
sum= (n+n+1) or 2n+1
difference=n-(n+1)(distribute the negative)=n-n-1 =-1 Make sure to use parenthesis here or the answer will not come out right.
so the product of the sum and difference is -1(2n+1) or -2n-1
"plus 8" to this and you get -2n-1+8 or -2n+7
we are told this is equal to the sum of their squares.
their squares will be n^2 and (n+1)^2
so we have -2n+7={n^2)+(n+1)^2
(n+1)^2 is the same as (n+1)(n+1)
Foil to multiply and you get n^2+n+n+1 or n^2+2n+1
-2n+7=n^2+n^2+2n+1 or -2n+7=2n^2+2n+1
get everything on one side and equal to zero.
0=2n^2+4n-6
factor out a 2 and you get 0=2(n^2+2n-3)
divide both sides by 2
0=n^2+2n-3
factors of -3 that add and give you positive 2 are 3 and -1, so it factors to (x+3)(x-1)
x+3=0 or x-1=0
x=-3 or x=1
look at our question again and we see that we want two consecutive POSTIVE numbers, so x cannot be -3. so x=1. If x=1, then our 2 consecutive positive numbers are 1 and 2.
Check your answer:
(sum)(difference)+8=1^2+2^2
(3)(-1)+8=1+4
5=5
Do you want your math questions answered quickly and completely? Do you want to ask follow up questions like, “How do you do that?” or “Why did you do that?”? Email your questions to justin.sheppard.tech@hotmail.com and I will answer your questions within 12 hours of submission (usually within 2-3 hours of submission). I charge $2 per question and I will provide a .pdf file of the problem worked out. I will also offer to discuss the problem with you over the phone, through skype, or via an online whiteboard. Payments to be made securely through PayPal AFTER you are satisfied with the answer and services that I provide. This eliminates the possibility of being scammed (worst case scenario is that I provide you with your answer and you don’t pay me).
.
justin.sheppard.tech@hotmail.com